OpenStudy (anonymous):

a sprinter who averaged 30 ft/sec completed a race 1 sec before another sprinter who averaged 28 ft/sec. what was the distance of the race?

OpenStudy (anonymous):

420 ft

OpenStudy (anonymous):

x = distance travelled by any of the racer distance = speed x time x = 30 ft/sec (t1) x = 28ft/sec (t2) t1 = t2-1 30ft/sec (t2-1) = 28ft/sec(t2) 30t2 - 30 = 28t2 30t2-28t2=30 2t2 = 30 t2 = 15 secs x = 28(15) = 420 ft